# Decision Synthesis in Monetary Policy

### v1

@@ Acknowledgements
Research of Emily Tallman was partially supported by the U.S. National Science Foundation through NSF Graduate Research Fellowship Program grant DGE 2139754. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. Further, the views expressed in this paper are solely those of the authors and may differ from the official views of the Bank of Canada; no responsibility for the views expressed in this paper should be attributed to the Bank of Canada.
@@
The vector ${\bm{\tau}}(\mathbf{x})$ is implicitly defined to satisfy $E_{f}[\mathbf{s}_{j}(\mathbf{y},\mathbf{x}_{j})|\mathbf{x}]=\mathbf{m}_{f}$ where $\mathbf{m}_{f}$ is the chosen target score and expectation is with respect to
@@
At any given $\mathbf{x},$ the integrals in equations (8,9) are approximated using Monte Carlo integration based on random samples from each of the $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j}).$ This gives the ingredients for numerical search over ${\bm{\tau}}(\mathbf{x})$ using derivative-free optimization methods such as those used in the current paper.
@@
The general theory for constraining on exact or uncertain linear constraints in multivariate normal forecast distributions (e.g. West and Harrison, 1997, sect. 16.3.2) underlies these results used in that paper and here in the specific case of VAR models.
@@
with the following terms: a is a $n\times 1$ vector of intercepts, $\mathbf{A}_{1},\dots,\mathbf{A}_{p}$ are the $n\times n$ VAR coefficient matrices, $\mathbf{A}_{0}$ is a full-rank contemporaneous impact matrix, $\mathbf{0}_{n}$ is a $n\times 1$ zero vector and $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
@@
where $\mathbf{R}$ is a chosen $r\times nh$ constant matrix with full row rank, $\mathbf{r}$ is an $r\times 1$ vector mean and ${\bm{\Omega}}$ is the corresponding $r\times r$ variance matrix related to the restrictions. This can be regarded as a set of uncertain constraints $\mathbf{r}$ when ${\bm{\Omega}}$ is non-zero, or exact constraints $\mathbf{R}\mathbf{y}_{T+1:T+h}=\mathbf{r}$ in the case ${\bm{\Omega}}=\mathbf{0}$.
Conditioning the forecast distribution of eqn. (12) on this constraint information yields the updated distribution below; details follow from West and Harrison (sect. 16.3.2 1997) and Chan et al. (2023). The updated distribution for the path is
@@
The special case of exact constraints $\mathbf{R}\mathbf{y}_{T+1:T+h}=\mathbf{r}$ has constraint uncertainty matrix ${\bm{\Omega}}=\mathbf{0}$ so that $\mathbf{M}^{*}$ reduces to $\mathbf{M}^{*}=\mathbf{M}-\mathbf{A}\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}\mathbf{A}^{\prime}.$ This is relevant in applications where it is justifiable to assume exact constraints. In our policy setting, the targeted constraint vector $\mathbf{r}$ is just that, a target policy path, for example, so that there will always be some level of uncertainty. The starting point is uncertainty represented by the variance matrix $V(\mathbf{R}\mathbf{y}_{T+1:T+h})=\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}.$ As argued in Antolin-Diaz et al. (2021), adopting ${\bm{\Omega}}=\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}$ represents a position that admits relevant and conservative levels of such uncertainty. This is used in the application of our paper as we constrain on candidate values of the policy path in considering conditional forecasts, hedged with uncertainty.
@@
The above development implicitly involves constraints imposed on all of the structural shocks of the model. In conditional forecasting exercises, it is often necessary to focus on only specific structural shocks set at restricted values (exact or uncertain), to obtain finer control. For example, a forecast conditional on an increasing policy rate path might result in an increase in inflation (i.e., the “price puzzle”) as the forecast is driven by reduced form shocks which are a mix of the structural shocks (i.e., demand, supply, monetary). For further discussion, see Antolin-Diaz et al. (2021). Due to these issues, we impose structural restrictions so that changes in the policy rate are driven by monetary policy shocks (i.e., decisions from the Central Bank) rather than as reactions to other shocks. This can be done by considering the restrictions
@@
where $\mathbf{R}_{0}$ is a selection matrix for observable restrictions, $\mathbf{r}_{0}$ is a vector of restrictions on observables, and $\Omega_{0}$ is a matrix of covariance restrictions on observables. Thus, by setting $\mathbf{R},\mathbf{r}$ and ${\bm{\Omega}}$ as indicated in eqn. (15), the structural restrictions can be regarded as a specific case of  eqn. (13). In our applications, we set $\mathbf{r}_{0}=\mathbf{x}$ and $\mathbf{W}=\mathbf{0}$ for all ${\bm{\varepsilon}}_{T+1:T+h}$ except those associated with the monetary policy shock– those are kept unrestricted. This amounts to conditioning on a proposed decision vector $\mathbf{x}$ and assuming it is driven only by monetary policy shocks. Other conditioning assumptions are possible. For example, in addition to assuming the driving shock is monetary policy one could assume a vector of positive/negative demand shocks representing different future scenarios.

### v2

@@ Acknowledgements
Research of Emily Tallman was partially supported by the US National Science Foundation through NSF Graduate Research Fellowship Program grant DGE 2139754. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. Further, the views expressed in this paper are solely those of the authors and may differ from the official views of the Bank of Canada. No responsibility for the views expressed in this paper should be attributed to the Bank of Canada.
@@
The vector ${\bm{\tau}}(\mathbf{x})$ is implicitly defined to satisfy $E_{f}[\mathbf{s}_{j}(\mathbf{y},\mathbf{x}_{j})|\mathbf{x}]=\mathbf{m}_{f}$, where $\mathbf{m}_{f}$ is the chosen target score and expectation is with respect to
@@
At any given $\mathbf{x},$ the integrals in equations (10,11) are approximated using Monte Carlo integration based on random samples from each of the $p_{j}(\mathbf{y}|\mathbf{x},{\mathcal{M}}_{j}).$ This gives the ingredients for numerical search over ${\bm{\tau}}(\mathbf{x})$ using derivative-free optimization methods such as those used in the current paper.
@@
The general theory for constraining on exact or uncertain linear constraints in multivariate normal forecast distributions (e.g., West and Harrison, 1997, sect. 16.3.2) underlies these results used in that paper and here in the specific case of VAR models.
@@
with the following terms: a is a $n\times 1$ vector of intercepts, $\mathbf{A}_{1},\dots,\mathbf{A}_{p}$ are the $n\times n$ VAR coefficient matrices, $\mathbf{A}_{0}$ is a full-rank contemporaneous impact matrix, $\mathbf{0}_{n}$ is a $n\times 1$ zero vector, and $\mathbf{I}_{n}$ is the $n\times n$ identity matrix.
@@
where $\mathbf{R}$ is a chosen $r\times nh$ constant matrix with full row rank, $\mathbf{r}$ is an $r\times 1$ vector mean, and ${\bm{\Omega}}$ is the corresponding $r\times r$ variance matrix related to the restrictions. This can be regarded as a set of uncertain constraints $\mathbf{r}$ when ${\bm{\Omega}}$ is non-zero, or exact constraints $\mathbf{R}\mathbf{y}_{T+1:T+h}=\mathbf{r}$ in the case ${\bm{\Omega}}=\mathbf{0}$.
Conditioning the forecast distribution of eqn. (14) on this constraint information yields the updated distribution that follows; details follow from West and Harrison (sect. 16.3.2 1997) and Chan et al. (2025). The updated distribution for the path is
@@
The special case of exact constraints $\mathbf{R}\mathbf{y}_{T+1:T+h}=\mathbf{r}$ has constraint uncertainty matrix ${\bm{\Omega}}=\mathbf{0}$ so that $\mathbf{M}^{*}$ reduces to $\mathbf{M}^{*}=\mathbf{M}-\mathbf{A}\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}\mathbf{A}^{\prime}.$ This is relevant in applications where it is justifiable to assume exact constraints. In our policy setting, the targeted constraint vector $\mathbf{r}$ is just that, a target policy path, for example, so there will always be some level of uncertainty. The starting point is uncertainty represented by the variance matrix $V(\mathbf{R}\mathbf{y}_{T+1:T+h})=\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}.$ As argued in Antolin-Diaz et al. (2021), adopting ${\bm{\Omega}}=\mathbf{R}\mathbf{M}\mathbf{R}^{\prime}$ represents a position that admits relevant and conservative levels of such uncertainty. This is used in the application of our paper as we constrain on candidate values of the policy path in considering conditional forecasts, hedged with uncertainty.
@@
The development in the previous section implicitly involves constraints imposed on all of the structural shocks of the model. In conditional forecasting exercises, it is often necessary to focus on only specific structural shocks, set at restricted values (exact or uncertain), to obtain finer control. For example, a forecast conditional on an increasing policy rate path might result in an increase in inflation (i.e., the “price puzzle”) as the forecast is driven by reduced form shocks that are a mix of the structural shocks (i.e., demand, supply, monetary). For further discussion, see Antolin-Diaz et al. (2021). Due to these issues, we impose structural restrictions so that changes in the policy rate are driven by monetary policy shocks (i.e., decisions from the Central Bank) rather than as reactions to other shocks. This can be done by considering the restrictions
@@
where $\mathbf{R}_{0}$ is a selection matrix for observable restrictions, $\mathbf{r}_{0}$ is a vector of restrictions on observables, and $\Omega_{0}$ is a matrix of covariance restrictions on observables. Thus, by setting $\mathbf{R},\mathbf{r}$, and ${\bm{\Omega}}$ as indicated in eqn. (17), the structural restrictions can be regarded as a specific case of  eqn. (15). In our applications, we set $\mathbf{r}_{0}=\mathbf{x}$ and $\mathbf{W}=\mathbf{0}$ for all ${\bm{\varepsilon}}_{T+1:T+h}$ except those associated with the monetary policy shock– those are kept unrestricted. This amounts to conditioning on a proposed decision vector $\mathbf{x}$ and assuming it is driven only by monetary policy shocks. Other conditioning assumptions are possible. For example, in addition to assuming the driving shock is monetary policy, we could assume a vector of positive/negative demand shocks representing different future scenarios.